OpenStudy (anonymous):

factor : a. x^3 - 196: b. x^2 +36: c. 15x^2-14x-8:

OpenStudy (anonymous):

b would involve imaginary numbers

OpenStudy (anonymous):

c would also since i come up with no rational factors

OpenStudy (anonymous):

not c i mean a

OpenStudy (anonymous):

can not factorable be an answer?

OpenStudy (anonymous):

yes but since b is involve with "i" how i do factor it?

OpenStudy (anonymous):

ok well a is not factorable

OpenStudy (anonymous):

imaginary numbers with all you would have to know is i^2 = -1

OpenStudy (anonymous):

so what do you suppse (x + 6I)(x - 6i) equals?

OpenStudy (anonymous):

ehh let me write in equation format \[(x + 6i)(x - 6i)\]

OpenStudy (anonymous):

make sense with the imaginary numbers?

OpenStudy (anonymous):

since FOIL would get you x^2 -6xi +6xi -36i^2 and i^2 = -1 so x^2 + 36

OpenStudy (anonymous):

yes makes sense :) and for A why would that be no rational numbers ?

OpenStudy (anonymous):

it isn't the fact that it isnt rational just 196 is not a perfect cube so there is not a way to factor it without more terms

OpenStudy (anonymous):

and c is (5x + 2)(3x - 4)

OpenStudy (anonymous):

ohh okay i got it lol

OpenStudy (anonymous):

yea factoring can be complicated at times but an easier example of b would be if x^2 - 36 since the perfect square factoring rule comes into play where (a^2 - b^2) = (a+b)(a-b)